a labourer is engaged for 30 days on the condition the he will be paid $80 for each day he works and he will be fine $10 for each day he is absent.if he finaly paid $1860. how many days did he come to work
Answer by ankor@dixie-net.com(22740)

a labourer is engaged for 30 days on the condition the he will be paid $80 for each day he works
he will be fined $10 for each day he is absent.
if he finally paid $1860. how many days did he come to work
:
let a = days absent
let w = days worked
:
a + w = 30
and
-10a + 80w = 1860
:
multiply the first equation by 10, add to the 2nd equation
10a + 10w = 300
-10a+ 80w =1860
------------------addition eliminates a, find w
0 + 90w = 2160
w = 2160/90
w = 24 days worked
:
:
Check this, 30-24 = 6 days absent
24*80 = 1920
6*10 = - 60
--------------
amt paid 1860